The proportion of right-handed students in a university population is 0.20, the probability of obtaining exactly 2 right handed people in a random sample of 4 is Select one: a. 0.154 b. 0.563 c. 0.432 d. 0.051 e. 0.381
Answer #1

Given that ,

p =0.20

1 - p = 1 - 0.20 = 0.80

(a)n =4

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X =2 ) = (4 C 2) * 0.202 * (0.80)4

=0.1536

probability =0.154